∫ (x)= ∫ (cosxy)dx + ∫ (siny²) dx


1
Expert's answer
2020-04-10T12:35:30-0400

"\int \cos{xy}dx+\int \sin{y^2}dx={\frac 1 y}\sin{xy}+x\sin{y^2}+C"